Upon graduating from college, you make an annual salary of $85,585. You set a goal to double it in the future. If your salary increases at an average annual rate of 8.62 percent, how long will it take to reach your goal?

Round the answer to two decimal places.
Answer #1
Starting Salary 85585
Interest 8.62%
Required salary level 171170
171170= 85585*(1+8.62%)^N
171170/85585= (1+8.62%)^N
2= 1.0862^N
Log 2= N log 1.0862
0.30103= N* 0.0359098
N= 0.30103/0.0359098
N=                          8.38 Years